OpenStudy (elonasushchik):

Point C is located at (1, 2), and point D is located at (−4, −2). Find the point that is the distance from point C to point D. (1 point) (−0.25, 1) (−1.5, 0) (−0.5, 1.5) (0.25, 2.25)

OpenStudy (elonasushchik):

@jabez177

OpenStudy (elonasushchik):

@sammixboo @SamsungFanBoy

OpenStudy (elonasushchik):

Which one do u think it is @ganeshie8

jabez177 (jabez177):

You can do these by yourself right ? I have no idea where the thing is...

OpenStudy (elonasushchik):

what thing lol ?! okayse

OpenStudy (elonasushchik):

@Najia2000

OpenStudy (elonasushchik):

@mathstudent55 @mathmate

OpenStudy (elonasushchik):

can someone guide me to the answer or at least ...

OpenStudy (mathmate):

Hint: Distance between two points = \(\sqrt{(x2-x1)^2+(y2-y1)^2}\)

OpenStudy (elonasushchik):

~23.3

OpenStudy (mathmate):

Sorry I gave you the wrong equation, because the question is not complete. I think the question should read: " Find the point that is \(equal\) distance from point C to point D. (1 point)" This means to find the mid-point between the two given points. The formula for mid-point is: mid-point=((x1+x2)/2, (y1+y2)/2) I think you will have no difficulty with that.

OpenStudy (elonasushchik):

So itsAor D right?

OpenStudy (mathmate):

I do not look at choices, so don't know which one is right. And I need you to show me how you got your answer if you want me to check it.

OpenStudy (mathmate):

I will be pleased to look at your answer, and how you got it.

OpenStudy (elonasushchik):

my equation right now is y=4/5x+6/5 4^2 + 5^2 = c^2 16 + 25 = 41 = c^2 so c = Sqrt(41)

OpenStudy (elonasushchik):

so square 41 is the distance between the two points

OpenStudy (elonasushchik):

Because the total rise is 4, so 1/4 of the rise will be just 1

OpenStudy (elonasushchik):

Since point C is (1,2) you want to add 1 to the y part then plug it in to your equation so plug 3 into the Y of your equation and solve (x,3) or 3 = 4/5x + 6/5 soo....umm

OpenStudy (elonasushchik):

so thats how I got D

OpenStudy (elonasushchik):

@mathmate u dere

OpenStudy (mathmate):

@ElonaSushchik First step is to confirm that my version of the question is correct, i.e. if YOUR question is asking for the position of the mid-point between C and D, i.e. the point which is equal distance from C & D.

OpenStudy (elonasushchik):

okay soo...

OpenStudy (elonasushchik):

was d correct?

OpenStudy (mathmate):

Can you first confirm the question? All four choices are wrong with the question as it is.

OpenStudy (elonasushchik):

yes it was correct Thank Yoou @mathmate

OpenStudy (mathmate):

The question is as posted?

OpenStudy (elonasushchik):

ya and I was correct

OpenStudy (mathmate):

Ok then I give up, because the question is asking for a distance, and all four choices are points. Perhaps you may want to sort this out with your teacher.